if a family drinks 3 1/2 gallons of milk in 5 days, how many gallons will they drink in 2 weeks?
Answer by tommyt3rd(5050)

3.5 : 5 = x : 14

3.5/5= x/14
x=14*3.5/5=9.8
9.8 gallons